1) Suppose x has a Poisson probability distribution with mean 4.84. Find standard deviation. 2)Assume that x has a Poisson probability distribution. Find P(x = 6) when population mean is 1.0. 3)Assume that x has a Poisson probability distribution. Find P(x < 3) when population mean is 4.5
Answer #1

1)

variance of poisson distribution = mean = 4.84

standard deviation = sqrt(variance)

= sqrt(4.84) = 2.2

2)

P(X) = e^-u (u^x)/x!

P(X=6) = e^-2.2 * (2.2)^6/6!

= 0.01745

3)

P(X<3) = P(X=0) + P(X=1) + P(X=2)

= e^-4.5 * (4.5^0)/1 + e^-4.5 * 4.5/1 + e^-4.5 * 4.5^2/2

= 0.1736